OpenStudy (anonymous):

Can somebody please explain to me how to solve inequalities ? like this one, -5 is greater than or equal to -5/9y

OpenStudy (shamil98):

\[\frac{ -5 }{ 9 }y \le -5\] multiply both side by the reciprocal of the fraction and flip the sign since you are multiplying by a negative number.

OpenStudy (shamil98):

so \[\frac{ -9 }{ 5 } (\frac{ -5 }{ 9 }) y \ge -5 \times \frac{ -9 }{ 5 }\]

OpenStudy (shamil98):

\[y \ge 9\]